OpenStudy (mony01):

find an equation of the tangent line to the curve at the given point x=1+sqrt(t), y=e^t^2 (2,e)

OpenStudy (dumbcow):

tangent line given point (x1, y1) \[y - y_1 = f'(x_1) (x - x_1)\] get y in terms of x take derivative

OpenStudy (mony01):

i took the derivative and this was my set up...is it right?\[\frac{ e ^{2t} }{ 2\sqrt{t} }\]

OpenStudy (dumbcow):

you need dy/dx , sub out the "t" variable first

OpenStudy (anonymous):

dy/dx = (dy/dt) / (dx/dt),

OpenStudy (dumbcow):

that works too

OpenStudy (anonymous):

\[dyd/x = \frac{2t e^{t^2}}{\frac{1}{2}t^{\frac{-1}{2}}}\]

OpenStudy (anonymous):

that's supposed to be *dy/dx*

OpenStudy (anonymous):

plug in t = 1

OpenStudy (anonymous):

y - e = m (x - 2) whatever the value of m is after you plug in 1 for t

OpenStudy (mony01):

is it 0 when you plug in 1 to the top equation? 2(1)e^(1)^2

OpenStudy (anonymous):

no it's 2e

OpenStudy (anonymous):

so m = 2e / (1/2) = 4e

OpenStudy (mony01):

is the answer y=4ex-9e?

OpenStudy (alekos):

y=4ex-7e